OpenStudy (anonymous):

Let f(x) = Find f(-2).who knows how to do this ?

hartnn (hartnn):

to find f(-2) put x = -2 in f(x)

OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). do you know how to this one because the other one wasnt the full equation